### v2

# Properties of Carbon-rich AGB Stars in the LMC and the Milky Way

Abstract
We present a comparative study of carbon-rich asymptotic giant branch (CAGB) stars in the Large Magellanic Cloud (LMC; 7347 stars) and the Milky Way (7163 stars) using infrared color–magnitude diagrams (CMDs), spectral energy distributions (SEDs), two-color diagrams (2CDs), and variability data. Observed SEDs are compared with theoretical models to characterize the central stars and their circumstellar dust envelopes and to estimate distances. For the LMC, a set of best-fitting CAGB models is derived by fitting observed SEDs with radiative transfer models, utilizing the galaxy’s well-established distance. For Galactic CAGB stars, where Gaia DR3 parallaxes are uncertain, we estimate distances by fitting observed SEDs with the CAGB models validated against LMC stars, and for Mira variables, from the period–magnitude relation calibrated with LMC Miras. A comparison of these approaches demonstrates that the SED-based distances are both reliable and practical for a large sample of Galactic CAGB stars. We find that CAGB stars in both galaxies show broadly similar infrared properties, although the LMC sample lacks stars with extremely thick dust envelopes.
Asymptotic giant branch stars (2100); Carbon stars (199); Long period variable stars (935); Circumstellar dust (236); Infrared astronomy (786); Radiative transfer (1335)
show]kwsuh@chungbuk.ac.kr
1 Introduction
The asymptotic giant branch (AGB) phase consists of two stages: the early AGB (E-AGB) and the thermally pulsing AGB (TP-AGB) (e.g., Iben & Renzini 1983). E-AGB stars are generally oxygen-rich (OAGB), whereas carbon-rich AGB (CAGB) stars are thought to evolve from intermediate-mass OAGB stars ($\approx$ 2 to 4.5 $M_{\odot}$ at solar metallicity, with the lower limit decreasing and the upper limit increasing at lower metallicity) once thermal pulses begin (e.g., Karakas 2014). More massive OAGB stars enter the TP-AGB phase without altering their surface chemistry. AGB stars are long-period variables (LPVs), characterized by large-amplitude pulsations and substantial mass-loss rates, which play a crucial role in driving envelope ejection and late stellar evolution (e.g., Suh 2014).
Although most carbon stars belong to the CAGB class, the category also encompasses other types. Classical (intrinsic, or type C-N) carbon stars are generally CAGB stars, whereas non-classical, or extrinsic, carbon stars are not on the AGB. The latter group includes CH stars, Barium stars, dwarf carbon stars, early R-type stars, and J-type stars (e.g., Green 2013; Suh 2024).
IR observations from Infrared Astronomical Satellite (IRAS), Midcourse Space Experiment (MSX), Infrared Space Observatory (ISO), AKARI, Two-Micron All-Sky Survey (2MASS), Wide-field Infrared Survey Explorer (WISE), and Spitzer have been instrumental in identifying CAGB stars and investigating their properties.
Gaia Data Release 3 (DR3) has yielded precise optical measurements for over one billion stars (Rimoldini et al. 2023). The distance determinations (Bailer-Jones et al. 2021) and extinction estimates (e.g., Lallement et al. 2022) derived from Gaia DR3 constitute essential inputs for constraining the absolute magnitudes of different stellar populations in the Milky Way.
Table 1: Sample of CAGB stars in the LMC and the Milky Way
Group Reference Total Number Selected
LMC-CAGB_SAGE Riebel et al. (2012) 9293 7308a
LMC-CAGB_SAGE-Sb SJc - 151
LMC-CAGB_KDMK Kontizas et al. (2001) 7760d 24+15e
LMC-CAGB combined - 7347
MW-CAGB_IC Suh (2024) 4909 4909
MW-CAGB_IC-S Suh (2024) - 91
MW-CAGB_NI Suh (2024) 2254 2254
MW-CAGB combined - 7163
aSee Section 2.1. bSubgroup of LMC-CAGB_SAGE; objects identified from the Spitzer IRS. cSloan et al. (2016); Jones et al. (2017). d3933 objects are duplicates of LMC-CAGB_SAGE objects. A large fraction of this sample could be extrinsic carbon stars. eMiras from OGLE-III and candidates for new Miras identified from WISE data (Suh 2025).
For Galactic CAGB stars, distance estimates based on Gaia DR3 parallaxes remain uncertain (e.g., Suh 2024). In this study, we use the well-established distance to the LMC to validate our models and methodology. Distances to Galactic CAGB stars are then derived by fitting their observed SEDs with a set of theoretical CAGB models calibrated against LMC CAGB stars at the known distance. For Galactic CAGB stars that are Mira variables, distances can also be estimated from the period–magnitude relation (PMR) established for LMC Miras.
We analyze the properties of CAGB stars in the LMC and the Milky Way. Section 2 presents catalogs of known CAGB stars in both galaxies. Section 3 describes radiative transfer models for circumstellar dust shells and the derivation of model parameters, constrained by LMC CAGB stars. In Section 4, observed spectral energy distributions (SEDs) are compared with model predictions to characterize the central stars and their dust envelopes and to estimate theoretical distances. Section 5 compares observed infrared two-color diagrams (2CDs) and color–magnitude diagrams (CMDs) with theoretical models. Section 6 examines the period–magnitude relation (PMR) of Mira-type CAGB stars. Section 7 summarizes the main results.
2 Sample Stars
We use catalogs of CAGB stars in the LMC and the Milky Way from the available literature (see Table 1). For all sample stars, we cross-identified counterparts from Gaia DR3, 2MASS, IRAS, MSX, ISO, AKARI, WISE, Spitzer, SIMBAD, and the American Association of Variable Star Observers (AAVSO; international variable star index; version 2025 June 29; Watson et al. 2025). All objects in the sample are identified by their unique SIMBAD identifiers, and their Gaia counterparts have been confirmed through SIMBAD.
2.1 CAGB stars in the LMC
[…]
The Optical Gravitational Lensing Experiment III (OGLE-III; Soszyński et al. 2009) reported 91,995 LPVs in the LMC, including 1663 classified as Mira variables.
Kontizas et al. (2001) compiled a catalog of 7760 carbon stars in the LMC, a large fraction of which may be extrinsic carbon stars. Among these, 3933 are duplicates of LMC-CAGB_SAGE objects. Of the remaining sources, 39 are likely additional CAGB stars, 24 are identified as Miras from OGLE-III, and 15 are strong candidates for new Miras based on recently analyzed WISE light curves (Suh 2025).
When we combine the two samples, there are 7347 CAGB stars in the LMC (LMC-CAGB objects; see Table 1). There are 1184 Mira variables from OGLE-III in the list of CAGB stars in the LMC (LMC-CAGB objects).
To calculate the absolute magnitudes for LMC-CAGB stars, we adopt a distance of 49.6 kpc to the LMC, a commonly used approximation based on the precise measurement of 49.59 $\pm$ 0.54 kpc by Pietrzyński et al. (2019).
2.2 CAGB stars in the Milky Way
Suh (2024) compiled a catalog of CAGB stars in the Milky Way, classified into two groups: those with IRAS counterparts (MW-CAGB_IC), identified from the IRAS source catalog and generally corresponding to brighter or more isolated stars, and those without IRAS counterparts (MW-CAGB_NI), identified from the AllWISE or Gaia DR3 catalogs and typically representing fainter stars or objects in crowded regions. Most MW-CAGB_IC sources also have entries in the AllWISE or Gaia DR3 catalogs. In this study, we adopt the catalog of Galactic CAGB stars from Suh (2024) (MW-CAGB objects; see Table 1). The MW-CAGB_IC-S subset denotes those MW-CAGB_IC objects for which infrared spectral data are available (see Section 4).
Table 2: Visual and IR bands
[…]
Bp[0.5] 0.511 3552 Gaia DR3 Rimoldini et al. (2023)
G[0.6] 0.622 3229 Gaia DR3 Rimoldini et al. (2023)
Rp[0.8] 0.777 2555 Gaia DR3 Rimoldini et al. (2023)
J[1.2] 1.235 1594 2MASS Cohen et al. (2003)
H[1.7] 1.662 1024 2MASS Cohen et al. (2003)
K[2.2] 2.159 666.7 2MASS Cohen et al. (2003)
W1[3.4] 3.35 306.681 WISE Jarrett et al. (2011)
W2[4.6] 4.60 170.663 WISE Jarrett et al. (2011)
S3[5.8] 5.731 115.0 Spitzer A
S4[8.0] 7.872 64.9 Spitzer A
MA[8.3] 8.28 58.49 MSX Egan et al. (2003)
AK[9] 9 56.262 AKARI Murakami et al. (2007)
IR[12] 12 28.3 IRAS Beichman et al. (1988)
W3[12]b 12 28.3 WISE Jarrett et al. (2011)
MC[12.1] 12.13 26.51 MSX Egan et al. (2003)
MD[14.7] 14.65 18.29 MSX Egan et al. (2003)
AK[18] 18 12.001 AKARI Murakami et al. (2007)
ME[21.3] 21.34 8.8 MSX Egan et al. (2003)
W4[22] 22.08 8.284 WISE Jarrett et al. (2011)
S5[24] 23.68 7.17 Spitzer B
IR[25] 25 6.73 IRAS Beichman et al. (1988)
IR[60] 60 1.19 IRAS Beichman et al. (1988)
AK[65] 65 0.965 AKARI Murakami et al. (2007)
AK[90] 90 0.6276 AKARI Murakami et al. (2007)
IR[100] 100 0.43 IRAS Beichman et al. (1988)
AK[140] 140 0.1859 AKARI Murakami et al. (2007)
AK[160] 160 0.1487 AKARI Murakami et al. (2007)
2.3 Observational Data
Table 2 presents the visual and IR bands used in this study for constructing SEDs, CMDs and 2CDs. The table includes the reference wavelength ($\lambda_{ref}$) and zero-magnitude flux (ZMF) for each band, which are essential for deriving theoretical models (see Section 3) and building SEDs from observational data. For a comprehensive discussion of IR photometric data and ZMF values, refer to Section 2.1 of Suh (2021).
For all sample stars, we cross-matched counterparts from Gaia DR3, 2MASS, IRAS, MSX, ISO, AKARI, WISE, Spitzer, SIMBAD, OGLE-III, and AAVSO by selecting the nearest sources within the angular resolution (beam size) of each telescope (see Suh 2021; Suh 2024). All objects are identified by their unique SIMBAD identifiers, and their Gaia counterparts were verified through SIMBAD.
IRAS, AKARI, and MSX data are available for most Galactic CAGB stars, but only for a small subset of those in the LMC. In contrast, Spitzer data cover the majority of LMC-CAGB stars through the SAGE program, but only a limited fraction of Galactic CAGB stars through the Galactic Legacy Infrared Mid-Plane Survey Extraordinaire (GLIMPSE; Churchwell et al. 2009), which does not include observations in the S5[24] band. By comparison, WISE and 2MASS data are available for most CAGB stars in both the LMC and the Milky Way.
This study incorporates only high-quality observational data across all wavelength bands for photometric analysis (quality better than 1 for IRAS, MSX, and AKARI data; quality better than U for WISE data).
2.4 Distances and extinction for Galactic CAGB stars
Recent parallax distance data from Bailer-Jones et al. (2021) (r_med_geo: the median of the geometric distance) derived from Gaia DR3, are crucial for determining the distances of many classes of stars in our Galaxy. Gaia DR3 provides readily available distances for a significant fraction of CAGB stars.
Gaia detectors operating in the visual bands cannot reliably detect CAGB stars with thick dust envelopes, often resulting in missing data for these objects. In addition, the Gaia counterparts provided by SIMBAD may be inaccurate for some objects, as it is often difficult to identify reliable counterparts using available photometric data and theoretical models, especially in crowded regions. Furthermore, CAGB stars frequently exhibit non-symmetric motions in their outer envelopes (e.g., Kim & Taam 2012), which can lead to inaccurate distance measurements in Gaia DR3. Therefore, distances derived from comparative analyses between observed and theoretical model SEDs may provide alternative or more reliable estimates (see Section 4).
Once the distance and position of each object were established, we obtained visual-band extinction values ($A_{V}$) from the database of Lallement et al. (2022). Extinction values for Gaia and 2MASS bands were then derived using the optical-to-MIR extinction law described by Wang & Chen (2019).
The derived distance and extinction data are employed to create a range of CMDs and IR 2CDs. In this study, extinction corrections are applied to the colors and magnitudes for the IR CMDs and 2CDs constructed from 2MASS band data (see Table 2).
3 Theoretical Dust Shell Models
To calculate theoretical model SEDs for CAGB stars, we use radiative transfer models for spherically symmetric dust shells surrounding central stars. The radiative transfer code RADMC-3D (http://www.ita.uni-heidelberg.de/~dullemond/software/radmc-3d/), following the approaches of Suh & Kwon (2013) and Suh (2024), is applied. Detailed descriptions of the theoretical models and their limitations can be found in Suh (2020).
[…]
We adopt a stellar blackbody temperature ($T_{*}$) between 2000 and 3000 K and a stellar luminosity ($L_{*}$) ranging from $5\times 10^{3}$ to $1\times 10^{4}$ $L_{\odot}$.
Table 3: 37 models for typical CAGB starsa
Modelb Dustc $\tau_{10}$ $T_{*}$ (K) $L_{*}$d
CA1 AMC 0.0001 3000 5.65
AMC 0.001 3000 5.65
AMC 0.005 3000 5.65
AMC+SiC 0.005 3000 5.65
AMC 0.007 2700 6.5
AMC+SiC 0.007 2700 6.5
CA2 AMC 0.01 2700 6.5
AMC+SiC 0.01 2700 6.5
AMC 0.02 2700 6.5
AMC+SiC 0.02 2700 6.5
AMC 0.03 2700 6.5
AMC+SiC 0.03 2700 6.5
AMC 0.05 2500 6.5
AMC+SiC 0.05 2500 6.5
AMC 0.07 2500 6.5
AMC+SiC 0.07 2500 6.5
CA3 AMC 0.1 2500 6.7
AMC+SiC 0.1 2500 6.7
AMC 0.2 2500 6.7
AMC+SiC 0.2 2500 6.7
AMC 0.3 2500 6.7
AMC+SiC 0.3 2500 6.7
AMC 0.4 2500 6.7
AMC+SiC 0.4 2500 6.7
CA4 AMC 0.5 2300 8
AMC+SiC 0.5 2300 8
AMC 0.6 2300 8
AMC 0.7 2300 8
AMC 0.8 2300 8
AMC 0.9 2300 8
CA5 AMC 1 2200 9
AMC 1.2 2200 9
AMC 1.5 2200 9
CA6 AMC 2 2000 10
AMC 2.5 2000 10
AMC 3 2000 10
CA7 AMC 4 2000 10
aSee Section 3 for details. For all models, $T_{c}$ = 1000 K. bSeven representative CAGB models were used to produce CMDs and 2CDs. cAMC: pure AMC; AMC+SIC: a simple mixture of AMC and SiC (20% by mass) dStellar luminosity in $10^{3}L_{\odot}$. Based on this luminosity, the scaled logarithmic distances can be sampled using a Gaussian Monte Carlo spread (GS) in $\log_{10}d$ ($\sigma=0.0798$ dex) derived from LMC-CAGB stars (see Figure 1), corresponding to a total luminosity variation of approximately $2.1\times$ from minimum to maximum (see Section 3.1).
3.1 Modeling LMC-CAGB stars
The absolute luminosity of CAGB stars in the Milky Way remains uncertain due to large errors in distance estimates. In contrast, the LMC, with its well-determined distance, provides an ideal setting for investigating the intrinsic luminosity of CAGB stars.
For the full sample of 7347 LMC-CAGB stars (Table 1), we tested an extensive grid of radiative transfer models covering a broad range of parameter combinations to identify the most suitable set of CAGB models. The observed SEDs were fitted using a least-squares optimization method (see Sections 4 and 4.1).
Figure 1: The top panel shows CAGB model SEDs for seven representative models (see Table 3). For LMC-CAGB stars, the shows the distance distribution (middle) and the relations between dust optical depth and luminosity (bottom) derived from CAGB model SEDs (Table 3). See Section 3.1 for details.
Figure 2: For LMC-CAGB stars, the top panels show the relations between optical depth and luminosity derived using two different methods based on CAGB models. In the top panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The bottom panels present IR CMDs for LMC-CAGB stars, with the brown arrow indicating an object likely to be an extrinsic carbon star (see Section 4.1). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$ = 0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 3.1 for details.
We began the procedure with the initial set of CAGB models for Galactic CAGB stars derived by Suh (2024) and applied them to the LMC-CAGB stars, refining and adjusting the model parameters. The sample was divided into five groups according to the dust optical depth ($\tau_{10}$), and for each group we determined the stellar temperature, spanning 2000–3000 K, that best reproduced the observed SED shapes of the LMC-CAGB stars. This process yielded five temperature groups: 3000, 2700, 2500, 2300, and 2200 K. Using the same five $\tau_{10}$ groups, we then computed SED-based scaling distance distributions for a range of stellar luminosities from $5\times 10^{3}$ to $1\times 10^{4}$ $L_{\odot}$.
We adopted five luminosity groups (5.65, 6.5, 6.7, 8, and 9 $\times 10^{3}$ $L_{\odot}$) that minimized the difference between the derived scaling distances and the known LMC distance as the best-fitting values. Each luminosity group corresponds to the same temperature group, except for the subset with $\tau_{10}$ between 0.05 and 0.07 in the 2500 K temperature group, for which the adopted luminosity is 6.5 $\times 10^{3}$ $L_{\odot}$ instead of 6.7 $\times 10^{3}$ $L_{\odot}$. In addition, we prepared a sixth group representing Galactic CAGB stars with very large dust optical depths ($\tau_{10}>1.5$). See Table 3 and the middle panel of Figure 1.
The LMC-CAGB stars lack objects with very large dust optical depths ($\tau_{10}>1.5$), whereas many such objects are present among Galactic CAGB stars (e.g., Suh 2024), for which a stellar temperature of 2000 K reproduces the observed SEDs well. For these heavily obscured objects, we extended the trend derived for lower optical depths in the LMC to a sixth group with a stellar temperature of 2000 K and a luminosity of $1.0\times 10^{4}$ $L_{\odot}$ (see Table 3).
Table 3 lists the final set of the 37 CAGB models that best reproduced the observed SED shapes of LMC-CAGB stars and minimized the differences between the scaling and true distances. The models are arranged in order of increasing $\tau_{10}$ and include the corresponding dust opacity and stellar parameters ($T_{*}$ and $L_{*}$). The top panel of Figure 1 presents SEDs for seven representative models.
The middle panel of Figure 1 shows the relation between distance and luminosity, both derived from the CAGB model SEDs (Table 3). Across the five luminosity groups, the derived distances are consistently in good agreement with the true distance of the LMC. In this diagram, the areas of the filled circular markers are proportional to the square root of the number of objects at each point.
The best-fitting CAGB model (from Table 3) and the corresponding scaling distance were derived for each LMC-CAGB star. Reliable results were obtained for 7205 of the 7347 LMC-CAGB stars (see Section 4). The bottom panel of Figure 1 shows the logarithmic distribution of the derived distances for the 7205 LMC-CAGB stars. The median distance (49.6 kpc) agrees well with the established value for the LMC, and the small dispersion in $\log_{10}d$ (0.0798 dex) demonstrates the consistency and reliability of the CAGB models (Table 3).
The upper panels of Figure 2 show the relations between optical depth and luminosity for the LMC-CAGB stars, derived using two different methods based on the CAGB model SEDs. The upper-left panel shows the relation obtained using the luminosities derived by the method described in the previous paragraphs (see Table 3; Figure 1). In contrast, the upper-right panel presents an alternative approach in which the luminosity was determined from the derived distance scaled to the true LMC distance of 49.6 kpc. This method reveals the detailed luminosity distribution for each $\tau_{10}$, which exhibits larger scatter. When the data from the five luminosity groups based on $\tau_{10}$ (as listed in Table 3) are combined, both methods yield similar results. In these diagrams, the areas of the filled circular markers are proportional to the square root of the number of objects at each point.
The bottom panels of Figure 2 present WISE–2MASS CMDs for LMC-CAGB stars, plotting the absolute magnitudes M(W3[12]) and M(K[2.2]) against the K[2.2]-W3[12] color. The observed absolute magnitudes, computed assuming a distance of 49.6 kpc, agree reasonably well with the CAGB models.
We find that M(W3[12]) traces the absolute luminosity reasonably well and can serve as a useful proxy for the luminosity function. In contrast, M(K[2.2]) exhibits fainter magnitudes for brighter CAGB stars with larger dust optical depths, because circumstellar dust absorbs shorter-wavelength light and re-emits it at longer wavelengths. Therefore, absolute magnitudes in NIR bands or at shorter wavelengths cannot be used as reliable proxies for the luminosity function for CAGB stars.
The dispersion in $\log_{10}d$ (0.0798 dex; corresponding to roughly 20 % in distance; right panel of Figure 1) for LMC-CAGB stars likely arises from three factors: (1) positional variations within the LMC, whose radius is about 10 % of its mean distance, implying a $\sim$5 % dispersion in distance; (2) the fact that a given $\tau_{10}$ value can correspond to multiple luminosities, as stars with different initial masses may exhibit different luminosities; and (3) intrinsic luminosity variations caused by the pulsations of CAGB stars. The observed dispersion in $\log_{10}d$ can therefore be regarded as representing the overall uncertainty in luminosity, since the contribution of the first factor is relatively minor. The measured dispersion in $\log_{10}d$ corresponds to a luminosity dispersion of 0.160 dex for the CAGB models (see Table 3).
Figure 3: The top panels show IR CMDs for Galactic CAGB stars. For Galactic CAGB stars, distances were derived from CAGB model SEDs, with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). For CAGB stars in the LMC and the Milky Way, the middle panels show the relations between optical depth and luminosity, and the bottom panels show the luminosity functions. In the middle panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The total number of objects is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 3.2 for details.
3.2 Application to Galactic CAGB stars
After confirming the reliability and uncertainties of the CAGB models for LMC-CAGB stars, we extended the analysis to the Milky Way sample (7163 MW-CAGB objects; Table 1), using the 37 CAGB models listed in Table 3 with observed SEDs (see Sections 4 and 4.2 for details).
The top panels of Figure 3 present the WISE–2MASS CMDs for Galactic CAGB stars, plotting M(W3[12]) and M(K[2.2]) against the K[2.2]-W3[12] color. Distances were derived by fitting the observed SEDs with the CAGB models (see Table 3; Section 4). To account for uncertainties in the luminosity of the CAGB models for Galactic CAGB stars, we applied Gaussian Monte Carlo scatter (GS) in distance ($\log_{10}d$) with $\sigma=\pm 0.0798$ dex, consistent with the dispersion measured for LMC-CAGB stars (see the right panel of Figure 1). CAGB stars in the LMC and the Milky Way exhibit very similar distributions on the CMDs (see Figures 2 and 3) and show good overall agreement with the CAGB models, displaying a comparable degree of spread.
The middle panels of Figure 3 show the relations between the observed color (K[2.2]-W3[12]) and the luminosity derived from the CAGB model SEDs (Table 3) for CAGB stars in the LMC and the Milky Way. Objects with higher derived luminosities exhibit redder observed colors, consistent with the characteristic behavior of the CAGB models, in which larger dust optical depths correspond to higher luminosities (see Table 3). The number distributions reveal both similarities and differences between the two galaxies. Although the overall trends are similar, Galactic CAGB stars display a broader range of colors than the LMC-CAGB stars. In particular, the LMC lacks objects with very red K[2.2]-W3[12] colors corresponding to large dust optical depths ($\tau_{10}>1.5$) at brighter luminosities.
The bottom panels of Figure 3 show the W3[12] band luminosity functions for CAGB stars in both galaxies. They exhibit similar overall shapes, each resembling a combination of two roughly Gaussian components: a main one with M(W3[12]) between 8-8 and 10.5-10.5 mag, and a secondary one comprising luminous sources with M(W3[12]) between 10.5-10.5 and 14-14 mag. Compared with the Milky Way, the LMC contains fewer of these luminous sources (M(W3[12]) <10.5<-10.5 mag), which correspond to “extreme carbon stars” representing the late stage of the CAGB phase (see Gruendl et al. 2008). Among Galactic CAGB stars, MW-CAGB_NI objects, which are generally fainter than MW-CAGB_IC objects (see Section 2.2), are rare at this bright end of the luminosity function.
3.3 Limitations of the Theoretical Models
The dust shell models adopted in this study do not account for gas-phase radiative processes (see Section 3). However, CAGB stars exhibit a variety of gas-phase emission and absorption features in the NIR and MIR bands arising from circumstellar molecules such as C2, C2H2, HCN, and CN (e.g., Le Bertre 2005). Consequently, the discrepancies between the models and observations are expected to be more pronounced at wavelengths where gas-phase radiation is significant.
Moreover, the spherical symmetry assumed in the dust shell models neglects the possibility of nonspherical dust geometries. As a result, CAGB stars with asymmetric dust envelopes can display markedly different SEDs that diverge substantially from the predictions of the theoretical models (e.g., Suh 2016).
4 SEDs and Distances of CAGB stars
Suh (2024) compared observed SEDs with theoretical models for a subset of Galactic CAGB stars. In this work, we extend that analysis to the full samples of CAGB stars in both the LMC and the Milky Way (see Table 1), using the radiative transfer models described in Section 3.
For constructing the observed SEDs, photometric data were collected from Gaia DR3, 2MASS, IRAS, MSX, AKARI, Spitzer, and AllWISE. For LMC-CAGB stars, Spitzer IRS (Jones et al. 2017) spectral data were included when available. For Galactic CAGB stars, IR spectra from the IRAS Low Resolution Spectrograph (LRS; Kwok et al. 1997) and ISO Short Wavelength Spectrometer (SWS; Sylvester et al. 1999) were used when accessible.
The observed SEDs of LMC-AGB stars were fitted using a least-squares optimization method, resulting in the 37 CAGB models (Table 3; see Section 3.1 and Figure 1) that best reproduced the observed SED shapes and minimized the differences between the scaling and true distances of LMC-CAGB stars. These LMC-validated CAGB models were then applied to the full sample of Galactic CAGB stars.
For the complete samples of CAGB stars in the LMC and the Milky Way, the best-fitting CAGB model ($\tau_{10}$) and the corresponding scaling distance, together with its uncertainty, were derived through least-squares optimization using the set of 37 CAGB models (Table 3). To ensure the reliability of the results, only fits with a coefficient of determination ($R^{2}$) exceeding 0.6 between the observed and model SEDs were retained. Applying this criterion, reliable best-fitting CAGB models and scaling distances were obtained for 7205 of the 7347 LMC-CAGB stars (98.1 %), 4884 of the 4909 MW-CAGB_IC stars (99.5 %), and 2238 of the 2254 MW-CAGB_NI stars (99.3 %).
Figure 4 compares the observed SEDs with CAGB models (see Table 3) for eight representative CAGB stars in the LMC (upper four panels) and the Milky Way (lower four panels), arranged in order of increasing dust optical depth ($\tau_{10}$). The comparisons show an overall similarity in the observed SEDs of CAGB stars in the LMC and the Milky Way.
[…]
4.1 SEDs of LMC-CAGB stars
The upper four panels of Figure 4 show the SEDs of four representative LMC-CAGB stars belonging to the subgroup LMC-CAGB_SAGE-S, for which infrared spectral data (Spitzer IRS) are available. The observed SEDs are reasonably well reproduced by the theoretical model SEDs (Table 3).
The two panels in Figures 5 compare the SEDs of two stars from the LMC-CAGB sample. The object in the left panel (MSX LMC 341) appears to be a typical CAGB star with a thick dust envelope. In contrast, the object in the right panel (2MASS J05454634-6732393) is the only source with a scaled distance greater than 100 kpc and does not resemble a typical CAGB star. Its SED shape is unusual, and the best-fitting CAGB model implies an extremely large distance ($\sim$131 kpc). If the object is instead located at the distance of the LMC (49.6 kpc), its absolute luminosity would be only $\sim$1200 $L_{\odot}$, about 6 times fainter than that of a typical CAGB star. This object is therefore more likely an extrinsic carbon star, which is thought to be a binary system consisting of a giant (or dwarf) star and a white dwarf that was once a CAGB star (see Suh 2024). The brown arrow in the CMDs of LMC-CAGB stars (see the bottom panels of Figure 2) marks the location of the object.
Figure 4: Observed and model SEDs for CAGB stars. The upper four panels show LMC stars, and the lower four show Milky Way stars. Each panel indicates the $\tau_{10}$ value and the corresponding scaling distance (bottom right). A suffix “s2” on $\tau_{10}$ denotes AMC+SiC dust opacity (20 % SiC). See Section 4 for details.
Figure 5: Observed and model SEDs for two LMC-CAGB stars with thick dust envelopes. The object in the right panel is likely an extrinsic carbon star (see Section 4.1).
Figure 6: Observed and model SEDs for Galactic CAGB stars with very thick dust envelopes ($\tau_{10}>1.5$). Such heavily obscured objects are not seen among LMC-CAGB stars (see Section 4.2).
4.2 SEDs of Galactic CAGB stars
The subgroup of MW-CAGB_IC objects (see Table 1) designated as MW-CAGB_IC-S consists of 91 well-studied CAGB stars with available infrared spectral data (IRAS LRS or ISO SWS). The lower four panels of Figures 4 show observed SEDs together with CAGB model SEDs for four representative Galactic CAGB stars from this subset. The SEDs of Galactic CAGB stars exhibit characteristics that are closely comparable to those of LMC-CAGB stars.
Figure 6 presents observed and model SEDs for four representative stars with very large dust optical depths ($\tau_{10}>1.5$), which are not seen among LMC-CAGB stars.
4.3 Distances of Galactic CAGB stars
For Galactic CAGB stars, the top panels of Figure 7 compare distances derived from Gaia DR3 parallax measurements with those obtained from theoretical CAGB model SEDs, shown separately for MW-CAGB_IC (left) and MW-CAGB_NI (right) stars. Although considerable scatter is present, the two distance estimates are generally consistent, with MW-CAGB_IC objects showing better overall agreement. The upper-left panel also highlights a comparison for the MW-CAGB_IC-S subgroup, a subset of CAGB_IC stars (see Section 4.2).
Of the 4909 MW-CAGB_IC objects (see the upper-left panel of Figure 7), 78.6% fall within a $\pm$0.23 dex agreement between the two distance estimates, as indicated by the lime dotted lines in the panel. This corresponds to distances agreeing within a factor of 1.7 (approximately -41% to +70% difference). When restricting the sample to the 114 objects with distances from the CAGB model SEDs within 900 pc, the agreement improves to 84.2%. Of the 2254 MW-CAGB_NI objects (see the upper-right panel of Figure 7), 55.9% fall within a $\pm$0.23 dex agreement between the two distance estimates.
The middle panels of Figure 7 compare distances derived from the period–magnitude relation (PMR) calibrated with LMC Miras (see Section 6; Equation 1) with those obtained from CAGB model SEDs for Mira-type CAGB stars. The two methods show strong consistency: among 1035 Mira-type MW-CAGB_IC stars (middle-left panel), 84.4 % agree within $\pm$0.23 dex, while for 114 Mira-type MW-CAGB_NI stars (middle-right panel), 71.1 % fall within the same range. Although both approaches have inherent limitations and uncertainties, SED fitting and the PMR provide the only practical means of deriving reliable distances from observed photometric and variability data for large samples of CAGB stars.
The lower-left panel of Figure 7 shows the WISE–2MASS CMD for Galactic CAGB stars, using distances derived from Gaia DR3 parallaxes. The diagram exhibits considerable scatter, with many absolute magnitudes lying outside the luminosity-function range for CAGB stars discussed in Section 3.2, and falling within the region occupied exclusively by extrinsic carbon stars, as identified by Suh (2024).
In contrast, the lower-right panel of Figure 7 presents the WISE–2MASS CMD for Galactic CAGB stars, using distances derived from theoretical CAGB model SEDs. The CMD exhibits significantly less scatter, with the resulting absolute magnitudes largely confined within the expected luminosity function limits. This indicates that distances derived from model SEDs are, in general, more reliable than Gaia DR3 parallax-based distances for a large sample of CAGB stars.
Among the three distance estimates for Galactic CAGB stars, model-based SED distances are found to be both reliable and practical for a large sample of Galactic CAGB stars. Gaia DR3 parallaxes have inherent limitations for AGB stars, and PMR-based distances apply only to Mira variables. In contrast, CAGB model SED distances can be derived whenever observed SEDs of sufficient quality and wavelength coverage are available.
Figure 7: For Galactic CAGB stars, the upper 4 panels show comparisons between distances derived from the CAGB model SEDs, Gaia DR3 parallax measurements, and the PMR obtained from Miras in the LMC. The bottom panels show IR CMDs for Galactic CAGB stars using distances from Gaia DR3 (left) and CAGB model SEDs (right). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The error bars represent the uncertainties in the derived distances. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 4.3 for details.
Figure 8: The top and lower-left panels show IR 2CDs for CAGB stars in the LMC and the Milky Way, while the lower-right panel compares fluxes in the W4[22] and S5[24] bands. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 5.
Figure 9: IR CMDs for CAGB stars in the LMC and the Milky Way. For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$=0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). For Galactic CAGB stars, distances are derived from CAGB model SEDs with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 5.
5 2CDs and CMDs in IR Bands
Srinivasan et al. (2011) and Groenewegen et al. (2020) modeled CAGB stars in the LMC using different approaches, presenting various 2CDs and CMDs and comparing observations with theoretical models. In this study, we extend this work by employing additional IR wavelength bands for a larger sample of CAGB stars in both the LMC and the Milky Way, allowing a more detailed comparison between observations and theoretical models based on more sophisticated schemes.
The upper two panels of Figure 8 show WISE–2MASS 2CDs of CAGB stars in the LMC and the Milky Way, using W3[12]–W4[22] versus K[2.2]W3[12]. In the upper-left region of the WISE–2MASS 2CD (see the upper panels of Figure 8), a subset of LMC-CAGB stars deviates significantly from the theoretical models with $T_{c}=1000$ K. In contrast, only a few Galactic CAGB stars show similar behavior. A plausible explanation is that the W4[22] band sensitivity is insufficient to reliably detect faint CAGB stars in the LMC.
[…]
The overall distributions of CAGB stars on the 2CDs in Figure 8 are broadly similar between the two galaxies, except that the LMC sample lacks objects with very large dust optical depths ($\tau_{10}>1.5$).
A comparable patten appeared in the WISE–2MASS CMDs (Figures 2 and 3), which compares CAGB stars in the LMC and the Milky Way, plotting M(W3[12]) and M(K[2.2]) against K[2.2]-W3[12]. Likewise, Figure 9 presents Spitzer–WISE–2MASS CMDs extending to longer wavelengths (W4[22] and S5[24]). The left panel shows M(S5[24]) versus K[2.2]-W3[12] for LMC-CAGB stars, while the right panel displays M(W4[22]) versus K[2.2]-W3[12] for Galactic CAGB stars.
[…]
Galactic CAGB stars also show brighter absolute magnitudes at the blue ends on the CMDs, which may reflect the LMC’s lower metallicity. At lower metallicity, the mass range of LMC-CAGB stars is expected to have a smaller lower limit and a larger upper limit compared to the Milky Way (see Karakas 2014). Consequently, Galactic CAGB stars are likely more massive and thus brighter than their LMC counterparts at the blue ends of the CMDs.
6 IR properties of Mira variables
AGB stars exhibit long-period, large-amplitude pulsations. It is widely accepted that more evolved (or more massive) AGB stars tend to show stronger pulsation amplitudes, longer periods, and enhanced mass-loss rates (e.g., De Beck et al. 2010).
Studying the period–magnitude relation (PMR) of Mira variables is more straightforward in the LMC than in our Galaxy, as their distances are relatively uniform. Miras in the LMC consistently exhibit a well-defined single-sequence PMR that is clearer in the MIR bands than at shorter wavelengths (e.g., Soszyński et al. 2009; Suh 2020; Iwanek et al. 2021).
To investigate the variability of CAGB stars in the WISE W1[3.4] and W2[4.6] bands over the past 16 years, one can use multi-epoch photometry from AllWISE (2009–2010) and the final data release of the Near-Earth Object WISE Reactivation (NEOWISE-R) mission (Mainzer et al. 2014). The latter provides observations from 21 epochs, typically two per year between 2014 and 2023, plus one in 2024. Such analyses have been carried out for Galactic CAGB stars by Suh (2024) and for LMC-CAGB stars by Suh (2025).
For LMC-CAGB stars, Suh (2025) considered both OGLE-III Miras and newly identified Mira candidates from WISE data, deriving the PMR in the W3[12] band. After correcting a few misidentified AllWISE counterparts in the sample, we derived a slightly revised PMR (see the top panel of Figure 10). From a second-order fit to the 1401 objects, the resulting relation is
$\mathrm{M(W3[12])}=-3.95(\log_{10}P)^{2}+11.7(\log_{10}P)-14.3\pm 0.65,$ (1)
where $P$ is the pulsation period in days.
Figure 10: PMRs for Mira-type CAGB stars. The top panel shows the PMR for LMC-CAGB objects and the bottom panel shows the PMR for MW-CAGB objects. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 6.
Figure 11: The upper two panels compare distances derived from the LMC-based PMR with those from Gaia DR3 for Mira-type Galactic CAGB stars. The bottom panel compares distances derived from the LMC-based PMR with those obtained from CAGB model SEDs for Mira-type LMC-CAGB stars. The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 6.
The bottom panel of Figure 10 presents the PMR for Galactic CAGB stars, combining known Mira variables from AAVSO with new Mira candidates identified from WISE data by Suh (2024), using distances derived from CAGB model SEDs. For the WISE-based candidates, only stars with periods longer than 400 days were included. Although the Galactic PMR exhibits greater scatter than that of the LMC, it broadly follows a similar trend.
For Galactic CAGB stars identified as Miras from AAVSO, the upper two panels of Figure 11 compare the alternative distances derived from the LMC-based PMR (Equation 1) with the Gaia DR3 distances. The vertical error bars for the PMR distances reflect the uncertainties in the LMC-based PMR. Among the 1006 Mira type MW-CAGB_IC objects shown in the top panel of Figure 11, 64.0% fall within a $\pm$0.23 dex agreement between the two distance estimates. For the 113 Mira type MW-CAGB_NI objects shown in the middle panel of Figure 11, 32.7% lie within this range of agreement. These results indicate that overall trends and correlations are consistent with those obtained from the CAGB model SEDs (see Figure 7).
For Mira-type LMC-CAGB stars, the bottom panel of Figure 11 compares distances derived from the LMC-based PMR (Equation 1) with those obtained from the CAGB model SEDs. Both methods show a similar correlation with the true distance, although the PMR-based distances exhibit a larger dispersion. These results further support the reliability of the distance estimates derived from the CAGB model SEDs.
7 summary
We have investigated the properties of CAGB stars in the LMC and the Milky Way, using samples of 7347 stars in the LMC and 7163 stars in the Milky Way. For all sample stars, we cross-identified counterparts from Gaia DR3, 2MASS, IRAS, MSX, ISO, AKARI, WISE, Spitzer, SIMBAD, and the AAVSO. Observed SEDs are compared with theoretical models to characterize the central stars and their circumstellar dust envelopes and to estimate distances.
For LMC-CAGB stars, we obtained a set of 37 CAGB models that best reproduced the shapes of the observed SEDs and minimized the differences between the scaling and true distances. The resulting distance distribution, with a median of 49.6 kpc and a small dispersion, confirms the consistency and reliability of the adopted CAGB models.
For Galactic CAGB stars, distance estimates from Gaia DR3 parallaxes remain uncertain. To address this, we derived distances by fitting observed SEDs with the CAGB models validated against LMC-CAGB stars. In addition, for Mira type Galactic CAGB stars, we determined distances using the period–magnitude relation (PMR) calibrated from LMC Miras.
Among the three distance estimates for Galactic CAGB stars, model-based SED distances were found to be both reliable and practical for a large sample of Galactic CAGB stars. Gaia DR3 parallaxes have inherent limitations for AGB stars, and PMR-based distances apply only to Mira variables. In contrast, CAGB model SED distances can be derived whenever observed SEDs of sufficient quality and wavelength coverage are available.
Overall, CAGB stars in the LMC and the Milky Way exhibit broadly similar properties in IR CMDs, SEDs, and 2CDs, but an important difference emerges: the LMC sample lacks stars with very large dust optical depths. Such heavily obscured CAGB stars are generally thought to represent the most massive or most highly evolved members of the population. Their relative scarcity in the LMC may therefore be a consequence of the galaxy’s lower rate of high-mass star formation compared with the Milky Way.
We have updated the catalog of Galactic CAGB stars from Suh (2024) by including three independent distance estimates for each object, derived in this work from Gaia DR3 parallaxes, CAGB model SED fitting, and the PMR for LMC Miras. The dataset for the updated catalog will be available on doi: https://doi.org/10.5281/zenodo.17262756.
I thank the anonymous reviewer for constructive comments and suggestions. This research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2022-NR075638). This research has made use of the SIMBAD database and VizieR catalogue access tool, operated at CDS, Strasbourg, France. This research has made use of the NASA/ IPAC Infrared Science Archive, which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.